### v1

@@ 2 Preliminaries
2.1 Notation
Let $\mathcal{X}$ denote a countable example space and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ denote a binary hypothesis class. Let $\mathcal{X}^{\star}$ denote the set of all finite subsets of $\mathcal{X}$. For any $h\in\mathcal{H}$, an enumeration of $\operatorname{supp}(h)$ is any infinite sequence $x_{1},x_{2},\dots$ such that $\bigcup_{i\in\mathbbm{N}}\{x_{i}\}=\operatorname{supp}(h)$. In other words, for every $x\in\operatorname{supp}(h)$, there exists an $i\in\mathbbm{N}$ such that $x_{i}=x$ and for every $i\in\mathbb{N}$, we have that $x_{i}\in\operatorname{supp}(h).$ For any $h\in\mathcal{H}$, a noisy enumeration of $\operatorname{supp}(h)$ is any infinite sequence $x_{1},x_{2},\dots$ such that for every $x\in\operatorname{supp}(h)$, there exists an $i\in\mathbb{N}$ such that $x_{i}=x$ and
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\not\in\operatorname{supp}(h)\}<\infty.$
For a finite sequence of examples $x_{1},\dots,x_{d}$ and $n\in\mathbb{N}\cup\{0\}$, define
$\mathcal{H}(x_{1},\dots,x_{d},n):=\{h\in\mathcal{H}:|\{x_{1},\dots,x_{d}\}\cap\operatorname{supp}(h)|\geq d-n\}.$
For any class $\mathcal{H}$, define its induced closure operator at noise level $n$ as $\langle\cdot\rangle_{\mathcal{H},n}$ such that
$\langle x_{1},\dots,x_{d}\rangle_{\mathcal{H},n}:=\begin{cases}\bigcap_{h\in\mathcal{H}(x_{1},\dots,x_{d},n)}\operatorname{supp}(h),&\text{if $|\mathcal{H}(x_{1},\dots,x_{d},n)|\geq 1$}\\ \bot,&\text{if $|\mathcal{H}(x_{1},\dots,x_{d},n)|=0$}\end{cases}.$
@@ 2.1 Notation
Assumption 1 (Uniformly Unbounded Support (UUS)).
@@
The UUS assumption is mainly needed for bookkeeping purposes to prevent the adversary from exhausting all positive examples and thus making the task of generating unseen positive examples impossible. This is consistent with the assumption that each language is countable infinite in size in the work of Kleinberg and Mullainathan (2024).
Finally, we will let $[N]:=\{1,\dots,N\}$ and abbreviate a finite sequence $x_{1},\dots,x_{n}$ as $x_{1:n}.$
@@ 2.2.1 Noiseless Generation
In the noiseless setting, an adversary plays a game against a generator $\mathcal{G}$. Before the game begins, the adversary picks a hypothesis $h\in\mathcal{H}$ and a sequence of examples $x_{1},x_{2},\dots$ such that $\{x_{1},x_{2},\dots\}\subseteq\operatorname{supp}(h).$ The adversary reveals the examples of stream to $\mathcal{G}$ one at a time, and the goal of the generator is to eventually output new, unseen positive examples $\hat{x}_{t}\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{t}\}$
Depending on how one quantifies “eventually,” one can get various notions of noiseless generatability. For example, if one requires to generator to perfectly generate new unseen example after observing $d\in\mathbbm{N}$ examples regardless of the hypothesis or stream chosen by the adversary, this is called as “uniform generation.”
Definition 2.2 (Uniform Generatability (Li et al., 2024)).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is uniformly generatable, if there exists a generator $\mathcal{G}$ and $d^{\star}\in\mathbbm{N}$, such that for every $h\in\mathcal{H}$ and any sequence $x_{1},x_{2},\dots$ with $\{x_{1},x_{2},\dots\}\subseteq\operatorname{supp}(h)$, if there exists $t^{\star}\in\mathbbm{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
If the number of positive examples required before perfect generation can depend on the hypothesis chosen by the adversary, but not the stream, then this is “non-uniform generation.”
Definition 2.3 (Non-uniform Generatability (Li et al., 2024)).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is non-uniformly generatable if there exists a generator $\mathcal{G}$ such that for every $h\in\mathcal{H}$, there exists a $d^{\star}\in\mathbbm{N}$ such that for any sequence $x_{1},x_{2},\dots$ with $\{x_{1},x_{2},\dots\}\subseteq\operatorname{supp}(h)$, if there exists $t^{\star}\in\mathbbm{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Finally, if the number of positive examples before perfect generation can depend on both the hypothesis and the stream selected by the adversary, this is called “generation in the limit.”
Definition 2.4 (Generatability in the Limit (Li et al., 2024)).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is generatable in the limit if there exists a generator $\mathcal{G}$ such that for every $h\in\mathcal{H}$, and any enumeration $x_{1},x_{2},\dots$ of $\operatorname{supp}(h)$, there exists a $t^{\star}\in\mathbbm{N}$ such that $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Note that uniform generation implies non-uniform generation which implies generation in the limit.
@@ 2.2.2 Noisy Generation
In the noisy model, we consider the following game. Like before, the adversary hypothesis $h\in\mathcal{H}$, and a sequence of positive examples $z_{1},z_{2},\dots\subseteq\operatorname{supp}(h)$. But now, the picks a noise level $n^{\star}\in\mathbb{N}$, and inserts at most $n^{\star}$ negative examples in $z_{1},z_{2},\dots$ to obtain a noisy stream $x_{1},x_{2},\dots$. The adversary then presents the examples in the noisy stream to generator $\mathcal{G}$ one at a time. Without knowledge of the noise-level or the location of the negative examples, the goal of the generator is to eventually output new, positive examples $\hat{x}_{t}\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{t}\}$.
@@
Like the noiseless case, there are several definitions for noisy generatability based on how one quantifies “eventually.” The most extreme case is what we term “uniform noisy generatability.”
Definition 2.5 (Uniform Noisy Generatability).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is uniformly noisily generatable, if there exists a generator $\mathcal{G}$ and $d^{\star}\in\mathbbm{N}$, such that for every $h\in\mathcal{H}$ and any sequence $x_{1},x_{2},\dots$ with
Here, the generator must perfectly generate after seeing $d\in\mathbb{N}$ positive example, where $d$ can only depend on the class $\mathcal{H}$ itself, and thus is uniform over the noise-level $n^{\star}$, the hypothesis $h\in\mathcal{H}$, and the stream $x_{1},x_{2},\dots$ selected by the adversary. In general, as we show in Section 3.1, uniform noisy generatability is impossible even for simple classes with just two hypotheses. This motivates weakening the definition by allowing $d$ to depend on the noise-level. We call this setting “Noisy Uniform Generatability.”
Definition 2.6 (Noisy Uniform Generatability).
@@
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is noisily uniformly generatable, if there exists a generator $\mathcal{G}$ such that for every noise level $n^{\star}\in\mathbb{N}$, there exists a $d^{\star}\in\mathbbm{N}$, such that for every $h\in\mathcal{H}$ and any sequence $x_{1},x_{2},\dots$ with
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n^{\star},$
if there exists $t^{\star}\in\mathbbm{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Noisy uniform generatability does not suffer from the same hardness of uniform noisy generatability. In fact, in Section 3.2, we show that all finite classes are uniformly noisily generatable. Nevertheless, we can continue weakening the notion of noisy generatability, by allowing $d$ to also depend on the hypothesis chosen by the adversary.
Definition 2.7 (Noisy Non-uniform Generatability).
@@
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is noisily non-uniformly generatable, if there exists a generator $\mathcal{G}$, such that for every noise level $n^{\star}\in\mathbb{N}$ and any $h\in\mathcal{H}$ there exists $d^{\star}\in\mathbbm{N}$ such that for any sequence $x_{1},x_{2},\dots$ with
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n^{\star},$
@@
if there exists $t^{\star}\in\mathbbm{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
@@
The term “non-uniform” here is used to refer to the fact that the number of unique examples needed by the generator before perfect generation can depend on the selected $h\in\mathcal{H}$, and hence it is “non-uniform” over the hypothesis class, unlike the previous two definitions. Finally, the weakest from of noisy generation allows $d$ to depend on the noise-level, the hypothesis, and the stream selected by the adversary.
@@
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is noisily generatable in the limit, if there exists a generator $\mathcal{G}$, such that for every $h\in\mathcal{H}$ and any noisy enumeration $x_{1},x_{2},\dots$ of $\operatorname{supp}(h)$, there exists $t^{\star}\in\mathbbm{N}$ such that $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
@@
We highlight that in all of these definitions, the generator is not aware of the noise-level or the noisy examples picked by the adversary. However, for all but “Uniform Noisy Generatability,” we allow the number of positive examples needed before perfect generation to depend on the noise-level in the stream.
Remark 2.3.
The astute reader might notice that there is actually a fifth setting of noisy generatability that we did not cover. In this fifth setting, which we will call “Non-uniform Noisy Generatability,” the number of positive examples needed before perfect generation can depend on the hypothesis chosen by the adversary, but must still be uniform over the noise-level and the noisy example stream chosen by the adversary. However, as in the case of uniform noisy generatability, non-uniform noisy generatability is still too strong as there exists a class of just two hypotheses that is not non-uniformly noisily generatable. We refer the reader to Appendix A for more details.
2.3 Summary of Results for Noiseless Generation
In the noiseless setting, Kleinberg and Mullainathan (2024) initiated the study of generation in the limit by showing that all countable classes are generatable in the limit and that all finite classes are uniformly generatable. Following this result, Li et al. (2024) provided a complete characterization of uniform generation in terms of a new combinatorial parameter termed the Closure dimension.
Definition 2.9 (Closure dimension (Li et al., 2024)).
The Closure dimension of $\mathcal{H}$, denoted $\operatorname{C}(\mathcal{H})$, is the largest natural number $d\in\mathbbm{N}$ for which there exists distinct $x_{1},\dots,x_{d}\in\mathcal{X}$ such that $\langle x_{1},\dots,x_{d}\rangle_{\mathcal{H},0}\neq\bot$ and $|\langle x_{1},\dots,x_{d}\rangle_{\mathcal{H},0}|<\infty.$ If this is true for arbitrarily large $d\in\mathbbm{N}$, then we say that $\operatorname{C}(\mathcal{H})=\infty.$ On the other hand, if this is not true for $d=1$, we say that $\operatorname{C}(\mathcal{H})=0.$
Theorem 2.1 (Theorem 3.3. in Li et al. (2024)).
Theorem 2.2 (Theorem 3.5 in Li et al. (2024)).
A class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is non-uniformly generatable if and only if there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\dots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{n}$ and $\operatorname{C}(\mathcal{H}_{n})<\infty$ for every $n\in\mathbb{N}$.
As a corollary, Li et al. (2024) establish that all countable classes are actually non-uniformly generatable. Charikar and Pabbaraju (2024) also established this result independently. With regards to generatability in the limit, Kleinberg and Mullainathan (2024) prove that all countable classes are generatable in the limit. Li et al. (2024) provide an alternate sufficiency condition in terms of the Closure dimension.
Theorem 2.3 (Theorem 3.10 in Li et al. (2024)).
@@ 3 Towards Characterizations of Noisy Generation
3.1 Uniform Noisy Generatability
We start by providing a characterization of the harshest form of noisy generation – uniform noisy generatability.
Theorem 3.1 (Characterization of Uniform Noisy Generatability).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. The following statements are equivalent.
  • (i)
    $\mathcal{H}$ is uniformly noisily generatable.
  • (ii)
    $\Bigl{|}\bigcap_{h\in\mathcal{H}}\operatorname{supp}(h)\Bigl{|}=\infty.$
Theorem 3.1 is a hardness result – it shows that even finite classes with just two hypotheses may not be uniformly noisily generatable. In fact, one way to interpret Theorem 3.1 is that uniform noisy generation is only possible for trivial classes where the generator can perfectly generate without observing any examples from the adversary. Indeed, this is what condition (ii) in Theorem 3.1 implies, since if condition (ii) is true, the generator can simply compute $\bigcap_{h\in\mathcal{H}}\operatorname{supp}(h)$ and always play from this set.

### v2

@@ 2 Preliminaries
Let $\mathcal{X}$ denote a countable example space and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ denote a binary hypothesis class. Let $\mathcal{X}^{\star}$ denote the set of all finite subsets of $\mathcal{X}$. Let $[N]:=\{1,\dots,N\}$ and abbreviate a finite sequence $x_{1},\dots,x_{n}$ as $x_{1:n}.$ For any $h\in\mathcal{H}$, an enumeration of $\operatorname{supp}(h)$ is any infinite sequence $x_{1},x_{2},\dots$ such that $\bigcup_{i\in\mathbb{N}}\{x_{i}\}=\operatorname{supp}(h)$. In other words, for every $x\in\operatorname{supp}(h)$, there exists an $i\in\mathbb{N}$ such that $x_{i}=x$ and for every $i\in\mathbb{N}$, we have that $x_{i}\in\operatorname{supp}(h).$ For any $h\in\mathcal{H}$, a noisy enumeration of $\operatorname{supp}(h)$ is any infinite sequence $x_{1},x_{2},\dots$ such that for every $x\in\operatorname{supp}(h)$, there exists an $i\in\mathbb{N}$ such that $x_{i}=x$ and $\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\not\in\operatorname{supp}(h)\}<\infty.$ For a finite sequence of examples $x_{1},\dots,x_{d}$ and $n\in\mathbb{N}\cup\{0\}$, define $\mathcal{H}(x_{1:d};n):=\{h\in\mathcal{H}:|\{x_{1:d}\}\cap\operatorname{supp}(h)|\geq d-n\}.$ For any class $\mathcal{H}$, define its induced closure operator at noise-level $n$ as $\langle\cdot\rangle_{\mathcal{H},n}$ such that
$\langle x_{1:d}\rangle_{\mathcal{H},n}:=\begin{cases}\bigcap_{h\in\mathcal{H}(x_{1:d};n)}\operatorname{supp}(h),&\text{if $|\mathcal{H}(x_{1:d};n)|\geq 1$}\\ \bot,&\text{if $|\mathcal{H}(x_{1:d};n)|=0$}\end{cases}.$
@@
Assumption 2.1 (Uniformly Unbounded Support (UUS) (Li et al., 2024)).
@@
The UUS assumption is mainly needed for bookkeeping purposes to prevent the adversary from exhausting all positive examples and thus making the task of generating unseen positive examples impossible. This is consistent with the assumption that each language is countably infinite in size in the work of Kleinberg & Mullainathan (2024).
@@ 2.1.1 Noiseless Generation
In the noiseless setting, an adversary plays a game against a generator $\mathcal{G}$. Before the game begins, the adversary picks a hypothesis $h\in\mathcal{H}$ and a sequence of examples $x_{1},x_{2},\dots$ such that $\{x_{1},x_{2},\dots\}\subseteq\operatorname{supp}(h).$ The adversary reveals the examples as a stream to $\mathcal{G}$ one at a time, and the goal of the generator is to eventually output new, unseen positive examples $\hat{x}_{t}\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{t}\}$.
Depending on how one quantifies “eventually,” one can get various notions of noiseless generatability. For example, if one requires the generator to perfectly generate new unseen examples after observing $d\in\mathbb{N}$ examples regardless of the hypothesis or stream chosen by the adversary, this is called “uniform generation.” If the number of positive examples required before perfect generation can depend on the hypothesis chosen by the adversary, but not the stream, then this is “non-uniform generation.” Finally, if the number of positive examples before perfect generation can depend on both the hypothesis and the stream selected by the adversary, this is called “generation in the limit.” We refer the reader to Appendix A for complete definitions and Appendix B for a summary of results.
@@ 2.1.2 Noisy Generation
In the noisy model, we consider the following game. Like before, the adversary picks a hypothesis $h\in\mathcal{H}$, and a sequence of positive examples $z_{1},z_{2},\dots\in\operatorname{supp}(h)$. But now, the adversary picks a noise-level $n^{\star}\in\mathbb{N}$, and inserts at most $n^{\star}$ negative examples in $z_{1},z_{2},\dots$ to obtain a noisy stream $x_{1},x_{2},\dots$. The adversary then presents the examples in the noisy stream to generator $\mathcal{G}$ one at a time. Without knowledge of the noise-level or the location of the negative examples, the goal of the generator is to eventually output new, positive examples $\hat{x}_{t}\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{t}\}$.
@@
Like the noiseless case, there are several definitions for noisy generatability based on how one quantifies “eventually.” The most extreme case is what we term “uniform noise-independent generatability.”
Definition 2.4 (Uniform Noise-Independent Generatability).
@@
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is uniformly noise-independent generatable, if there exists a generator $\mathcal{G}$ and $d^{\star}\in\mathbb{N}$, such that for every $h\in\mathcal{H}$ and any sequence $x_{1},x_{2},\dots$ with $\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}<\infty$, if there exists $t^{\star}\in\mathbb{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Here, the generator must perfectly generate after seeing $d^{\star}\in\mathbb{N}$ examples, where $d^{\star}$ can only depend on the class $\mathcal{H}$ itself, and thus is uniform over the noise-level $n^{\star}$, the hypothesis $h\in\mathcal{H}$, and the stream $x_{1},x_{2},\dots$ selected by the adversary. Unsurprisingly, as we show in Section 3.1, uniform noise-independent generatability is impossible even for simple classes with just two hypotheses. In Appendix C, we show this is also the case if we measure sample complexity in terms of just the number of distinct positive examples (as opposed to all examples). These results motivate weakening the definition by allowing $d^{\star}$ to depend on the noise-level. We call this setting “Uniform Noise-dependent Generatability.”
Definition 2.5 (Uniform Noise-dependent Generatability).
@@
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is uniformly noise-dependent generatable, if there exists a generator $\mathcal{G}$ such that for every noise-level $n^{\star}\in\mathbb{N}$, there exists a $d^{\star}\in\mathbb{N}$, such that for every $h\in\mathcal{H}$ and any sequence $x_{1},x_{2},\dots$ with $\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n^{\star},$ if there exists $t^{\star}\in\mathbb{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Uniform noise-dependent generatability does not suffer from the same hardness of uniform noise-independent generatability. In fact, in Section 3.2, we show that all finite classes are uniformly noise-dependent generatable. Nevertheless, we can continue weakening the notion of noisy generatability, by allowing $d^{\star}$ to also depend on the hypothesis chosen by the adversary.
Definition 2.6 (Non-uniform Noise-dependent Generatability).
@@
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is non-uniformly noise-dependent generatable, if there exists a generator $\mathcal{G}$, such that for every noise-level $n^{\star}\in\mathbb{N}$ and any $h\in\mathcal{H}$ there exists $d^{\star}\in\mathbb{N}$ such that for any sequence $x_{1},x_{2},\dots$ with $\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n^{\star},$ if there exists $t^{\star}\in\mathbb{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
@@
The term “non-uniform” here is used to refer to the fact that the number of unique examples needed by the generator before perfect generation can depend on the selected $h\in\mathcal{H}$, and hence it is “non-uniform” over the hypothesis class, unlike the previous two definitions. Finally, the weakest form of noisy generation allows $d$ to depend on the noise-level, hypothesis, and stream selected by the adversary.
@@
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is noisily generatable in the limit, if there exists a generator $\mathcal{G}$, such that for every $h\in\mathcal{H}$ and any noisy enumeration $x_{1},x_{2},\dots$ of $\operatorname{supp}(h)$, there exists $t^{\star}\in\mathbb{N}$ such that $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
@@
The astute reader might notice that there is actually a fifth setting of noisy generatability that we did not cover. In this fifth setting, which we will call “Non-uniform Noise-independent Generatability,” the number of positive examples needed before perfect generation can depend on the hypothesis chosen by the adversary, but must still be uniform over the noise-level and the noisy example stream chosen by the adversary. However, as in the case of uniform noise-independent generatability, non-uniform noise-independent generatability is still too strong as there exists a class of just two hypotheses that is not non-uniformly noise-independent generatable. We refer the reader to Appendix D for more details.
@@ 3 Towards Characterizations of Noisy Generation
3.1 Uniform Noise-independent Generatability
We start by providing a characterization of the strongest form of noisy generation – uniform noise-independent generatability.
Theorem 3.1 (Characterization of Uniform Noise-independent Generatability).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is uniformly noise-independent generatable if and only if $\Bigl{|}\bigcap_{h\in\mathcal{H}}\operatorname{supp}(h)\Bigl{|}=\infty.$
Theorem 3.1 is a hardness result – it shows that even finite classes with just two hypotheses may not be uniformly noise-independent generatable. In fact, one way to interpret Theorem 3.1 is that uniform noise-independent generation is only possible for trivial classes where the generator can perfectly generate without observing any examples from the adversary. Indeed, this is what condition in Theorem 3.1 implies, since if this condition is true, the generator can simply compute $\bigcap_{h\in\mathcal{H}}\operatorname{supp}(h)$ and always play from this set. In Appendix C, we prove a similar statement even if we only measure the sample complexity in terms of the number of distinct positive examples. Since the sufficiency direction of Theorem 3.1 follows from this observation, we only prove the necessity direction.